OpenStudy (anonymous):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (anonymous):

With the length, do you know which side is which length? If so, and you have two sides and an angle, a right triangle, I believe you have enough information.

OpenStudy (anonymous):

Well I can make the lengths whatever I want to. I have to give an example on how to do it. Like, we can have side BA = 57

OpenStudy (anonymous):

and AC = 86

OpenStudy (anonymous):

for right triangle, a^2+b^2=c^2. c is the hypotenuse. angle B=90 deg, angles A+C=90 deg. use sin and cos identities. or tan and cot.

OpenStudy (anonymous):

how would i use tan and cot to figure it out?

OpenStudy (anonymous):

sorry, i take the class online and really dont understand

OpenStudy (phi):

are you studying geometry or trigonometry. Did you learn about the law of sines or the law of cosines?

OpenStudy (anonymous):

i am studying geometry

OpenStudy (phi):

You need to know trigonometry to find the two (non-right) angles, given that you know the sides. See http://www.khanacademy.org/math/trigonometry/v/basic-trigonometry On the other hand, if you know one of the (non-right) angles, you can find the other because all 3 angles must sum to 180. And if you know 2 of the sides, you can find the 3rd using pythagorean theorem.

OpenStudy (phi):

Though sometimes you learn about "special triangles" in geometry: 30-60-90 and 45-45-90 In those cases, you can figure out the angles knowing the sides.

OpenStudy (anonymous):

Yeah, I know how to get the third side. it would be 64.40. i used the Pythagorean theorem.

OpenStudy (phi):

Have you learned about 30-60-90 triangles?

OpenStudy (anonymous):

yeah

OpenStudy (phi):

If you are making up an example you could pick special lengths: you can pick a hypotenuse AC= 10 and side BC= 5. Because BC is 1/2 of AC you know angle CAB is 30º and side AB is sqrt(3)*5

OpenStudy (phi):

and angle ACB would be 60º

OpenStudy (anonymous):